Graphs of the two functions y=2x and y=(1/2)x

A. Intersect at a point of negative abscissa

B. Intersect at a point of positive abscissa

C. Don't intersect

D. Intersect at a point of zero abscissa


1
Expert's answer
2021-08-09T15:13:21-0400
"2^x=(1\/2)^x"




"4^x=1"

"x=0"

Graphs of the two functions y=2x and y=(1/2)x intersect at point "(0,1)."


D. Intersect at a point of zero abscissa.